### v1

# Denoising Mutual Knowledge Distillation in Bi-Directional Multiple Instance Learning

Abstract
Multiple Instance Learning is the predominant method for Whole Slide Image classification in digital pathology, enabling the use of slide-level labels to supervise model training. Although MIL eliminates the tedious fine-grained annotation process for supervised learning, whether it can learn accurate bag- and instance-level classifiers remains a question. To address the issue, instance-level classifiers and instance masks were incorporated to ground the prediction on supporting patches. These methods, while practically improving the performance of MIL methods, may potentially introduce noisy labels. We propose to bridge the gap between commonly used MIL and fully supervised learning by augmenting both the bag- and instance-level learning processes with pseudo-label correction capabilities elicited from weak to strong generalization techniques. The proposed algorithm improves the performance of dual-level MIL algorithms on both bag- and instance-level predictions. Experiments on public pathology datasets showcase the advantage of the proposed methods.
1 Introduction
Multiple Instance Learning (MIL) is a widely used framework in computational pathology, particularly for Whole Slide Image (WSI) classification [1, 2, 3, 4, 5]. Due to the extremely high resolution of the WSIs, processing them as a whole is computationally expensive. Meanwhile, obtaining pixel-level or patch-level annotations is both time-consuming and expensive, which limits the scalability of fully supervised learning approaches. In the MIL framework, each WSI is partitioned into a set of smaller image patches. These patches, regarded as instances, are aggregated into a bag that shares the corresponding slide-level label. The model is trained using only bag-level labels, while the individual patches (instances) remain unlabeled. The MIL framework alleviates the reliance on fine-grained annotations, making MIL especially suitable for WSI classification.
Existing MIL approaches can be broadly categorized into two types: bag-level methods and instance-level methods. Bag-level methods aggregate the features of all instances within a bag embedding to predict the bag label, without explicitly modeling the class of individual instances. Most deep learning-based MIL belong to the prior category, including the widely used attention-based MIL (ABMIL) [6] and CLAM [7]. Instance-level methods make a prediction at the instance level and then infer the bag label based on instance scores, such as DSMIL [8]. MIL methods with integrated bag and instance branches have also been proposed to utilize the strengths of both paradigms, including approaches like TransMIL [9] and DTFD-MIL [10].
Both bag-level and instance-level methods have been extensively studied, each has its own advantages and drawbacks. Bag-level methods maximize the utility of the ground truth bag labels, allowing them to excel in bag classification tasks. However, the bag embedding provably contains less information than the whole set of instance features [11]. As a direct consequence, the bag-level methods tend to focus on learning only easy instances [12] or spurious correlations [13] within a bag, limiting their generalization ability. Instance-level methods consider all instances individually. However, the lack of ground-truth supervision for individual instances makes them susceptible to noise in the instance pseudo labels during training. These drawbacks in both categories compromise the performance and robustness of MIL models.
To address the issues of not exploiting sufficient instances in bag-level methods, hard instance mining (HIM) modules are proposed to mask easy instances during training [12, 14]. HIM modules encourage the model to focus on hard instances during training, thereby enhancing its ability to generalize. The IBMIL model [13] and the CAMIL model [15] are proposed recently to avoid associating bag classifiers with spurious features by eliminating harmful confounders using causal inference. The CausalMIL model [16, 17] utilizes causal inference in MIL to reduce its sensitivity to data distribution, finding stable instances to guide bag-level classification.
Previous work has also investigated methods to improve the reliability of the instance-level pseudo-labels. The CIMIL model [18] used counterfactual inference and a hierarchical instance searching module to search for reliable instances and generate precise pseudolables. Optimal transport is also used to calibrate soft labels in a recent paper [11]. Qu et al. [19] use prototype learning to improve the reliability of pseudolable based on representation learned by the the so-called instance-level weakly supervised contrastive learning.
Although current improvement methods have demonstrated some effectiveness, they may also introduce biases from other sources into the model. HIM may inadvertently introduce noise into the bag label when the masks filter out all positive instances, as discussed in prior work [20]. Causal inference methods depend on the accurate modeling of confounders, without which the inferred effects may be invalid [21]. In addition, some MIL frameworks adopt complex mechanisms to identify representative instances within a bag, resulting in increased implementation difficulty and reduced interpretability [22]. The use of combinatorial or multi-stage algorithms further increases computational cost, making such methods less scalable in real-world applications [23].
Without external guidance such as instance level annotation, leveraging the model’s own capability to correct pseudo-label noise seems inevitable [24]. The weak to strong generalization phenomenon describes the ability of a student to outperform weak teachers in terms of generalization capability when it is pretrained to generate robust data representations [25]. Our denoising method is based on this simple idea of amplifying generalization power using only weak supervision.
In this paper, we propose a dual-level learning algorithm that can fully utilize weak supervision in a mutual distillation process which encompasses bag- and instance-level training. Our main contributions are summarized as follows:
  • We propose a dual-level MIL architecture that include both bag branch and instance branch. Compared with previous methods, the supervision signals in both branches are enhanced: (1) the soft labels are anchored to specific classes to provide improved supervision for instance branch; (2) the instance predictions are aggregated to assist the training of the bag branch as a supplement to bag labels. Besides, the training process of the two branches are scheduled to ensure that they are optimized at the same pace.
  • We use self-confident predictions as a distillation signal to enhance the generalization capability of the models in both branches. To this end, self-confidence losses are added to both branches. In the bag branch, the additional loss further strikes a balance between the bag label and the aggregated instance predictions. Experiments on both the CAMELYON16 dataset and the TCGA-NSCLC dataset show strong prediction performance of the proposed methods at the bag and instance level.
2 Related Work
Deep multiple instance learning (Deep MIL) extends the classical MIL framework by integrating neural networks to jointly learn instance-level and bag-level representations. Wang et al. [26] proposed mi-Net and MI-Net as basic deep learning-based instance-level and embedding-level pooling methods. Ilse et al. [6] introduced the Attention-Based MIL model (ABMIL), which employs a learnable attention mechanism to aggregate instance embeddings within each bag. More recent work has used transformers [9] and graph neural networks [27, 28, 29] to aggregate instance features.
Instance classifiers are incorporated in many MIL models to ensure that bag predictions are consistent with key instance predictions. An instance classifier is a component in MIL frameworks that assigns labels or scores to individual instances, enabling finer-grained modeling under weak supervision. The CLAM model [30] uses an instance classification head to ensure that the top $K$ attended patches share the same class with the bag. The DGMIL model [31] uses instance-level pseudo-labels to refine the feature distribution to separate positive and negative instance representation. Ma et al. [11] treated MIL as a semi-supervised task, training an instance classifier with pseudo labels and consistency regularization. Qu et al. [19] leveraged contrastive and prototype-based learning to enhance the discriminative ability of instance classifiers. Lin et al. [18] proposed a counterfactual inference framework named CIMIL that uses instance representation to generate more informative representations for MIL models.
Hard instance mining has been widely adopted to enhance model robustness and generalization in scenarios involving ambiguous or weakly informative samples. In object detection, Shrivastava et al. [32] proposed Online Hard Example Mining (OHEM), which prioritizes training on samples with the highest loss to accelerate convergence and improve detection accuracy. This approach laid the groundwork for subsequent applications of hard mining in other domains.
In the context of applying MIL for WSI classification, the WENO model [12] uses a hard positive instance mining (HPM) module to mask the most attended patches in bag classification training to force the model to learn hard instances. Tang et al. [14] introduced a masked hard instance mining (MHIM) mechanism that identifies challenging instances by enforcing instance-level consistency within a teacher-student framework. Their methods improves MIL performance by explicitly guiding the model to focus on informative but difficult regions in histopathology images, outperforming previous attention-based MIL approaches. Other work leveraging HIM includes HPA-MIL [33] and ACMIL [34].
Weak-to-Strong (W2S) generalization studies methods to supervise a stronger model using supervision from a weaker model. Burns et al. [35] introduced this setting and demonstrated it empirically through several tasks. Charikar et al. [36] provided a theoretical framework quantifying how strong models can benefit from weak labels. Lang et al. [37] proposed mechanisms explaining how label correction and coverage expansion support generalization beyond the supervisor.
From the application perspective, Guo et al. [38] showed that vision foundation models trained with weak supervision can outperform both their weak supervisors and fully supervised baselines. Wang et al. [39] utilize confident subset of pseudo-labels to guide the instruction tuning of LLMs. Cui et al. [40] used aggregated weak feedback to fine-tune large language models on text classification and generation tasks, achieving higher-quality outputs. Somerstep et al. [41] proposed a refinement-based transfer learning approach that enables stronger models to generalize beyond weak teachers in alignment tasks.
3 Method
In this section, we begin by introducing the formulation of MIL problems. Then we formally describe attention-based MIL methods used for WSI classification, based on which we propose a method that avoids fitting to noisy supervision signals in a bi-directional distillation framework.
3.1 Formulation of MIL problems
In the MIL framework, the dataset consists of a set of bags $X\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{X_{1},X_{2},\dots,X_{N}\}$, where each bag $X_{i},\forall i\in[N]$ is a collection of $n_{i}$ instances $X_{i}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{x_{i}^{j}\}_{j=1}^{n_{i}}$. For a bag $X_{i},i\in N$, we denote by $Y_{i}\in\{0,1\}$ its bag label. Let $S\in[N]$ be the training set. The training algorithm takes as input $\{(X_{i},Y_{i})\}_{i\in S}$. The instance labels $\{y_{i}^{j},j\in[n_{i}]\},y_{i}^{j}\in\{0,1\}$ are only available in the testing phase. Under the standard MIL assumption, the relationship between the bag label $Y_{i}$ and all its instance labels is described as follows:
$Y_{i}=\left\{\begin{array}[]{ll}0&\text{if }\sum_{j}y_{i}^{j}=0\\ 1&\text{otherwise,}\end{array}\right.$ (3)
where all instances in negative bags are given negative instance labels, and at least one instance in a positive bag is designated with a positive label.
During the training phase, an MIL algorithm uses a set of $(X_{i},Y_{i})$ pairs to learn a bag-level classifier $f_{\text{bag}}:X\to\{0,1\}$ and possibly an instance-level classifier $f_{\text{inst}}:\mathcal{X}\to\{0,1\}$, where $\mathcal{X}$ consists of instances from all bags. In this paper, the proposed MIL algorithm aims to learn both bag- and instance-level classifiers with only bag-level labels provided in the training set.
3.2 MIL in WSI Classification
In a typical WSI classification workflow, a WSI is first partitioned into a set of non-overlapping patches, each of which is treated as an instance within a bag. These patches are individually passed through an instance encoder $\mathcal{E}$ to extract instance-level features. The resulting features are then aggregated by a feature aggregation module $g$ to obtain a holistic bag-level representation. This aggregated representation is subsequently processed by a classification head $\varphi$ to predict the bag-level label. The entire model is trained end-to-end using the bag-level labels as supervision.
A line of widely used MIL methods utilizes attention-based aggregation, which assigns a learnable attention score to each instance based on its relevance to the classification task, and then combines them via a weighted sum to produce the final bag representation. By enabling the model to selectively focus on the most informative regions, attention mechanisms significantly enhance both interpretability and performance in MIL-based WSI classification models.
Formally, given a bag $X_{i}=\{x_{i}^{j}\}_{j=1}^{n_{i}}$, each instance $x_{i}^{j}$ is encoded into a feature vector $h_{i}^{j}=\mathcal{E}(x_{i}^{j})\in\mathbb{R}^{d}$. An attention mechanism then computes a scalar attention score $a_{i,j}\in[0,1]$ for each instance $x_{i}^{j}$. In the ABMIL model, $a_{i,j}$ is defined as $a_{i}^{j}=w^{\top}\tanh{(V(h_{i}^{j})^{T})}$ ), where $V\in\mathbb{R}^{l\times d}$ and $w\in\mathbb{R}^{l}$ are learnable parameters. The vector $a_{i}=[a_{i}^{1},\dots,a_{i}^{n_{i}}]$ is then normalized by using a softmax function $\alpha_{i}^{j}=\nicefrac{{\text{exp}({a_{i}^{j}})}}{{\sum_{i=1}^{n_{i}}\text{exp}({a_{i}^{j}})}}$. The bag-level representation is obtained as a weighted average of instance features: $H_{i}=\sum_{j=1}^{n_{i}}\alpha_{i}^{j}\cdot\mathcal{E}(x_{i}^{j})$. This representation is then passed to a classifier $\varphi$ to predict the bag label.
3.3 Overview of the Model
Figure 1 illustrates the overall architecture of the proposed model, which is composed of two interconnected components: a bag-level branch and an instance-level branch. Notations in Figure 1 are explained in the remainder of the section.
Figure 1: Framework overview of the proposed method.
The bag-level branch is primarily responsible for learning a robust bag-level classifier using supervision provided by bag-level labels $Y_{i}$. It also generates attention scores $a_{i}^{j}$ (and normalized scores $\alpha_{i}^{j}$) for each instance $x_{i}^{j}$ within the bag $X_{i}$. These attention scores not only indicate the relative importance of individual instances to the bag classification but also serve as weak supervisory signals to guide the optimization of the instance-level branch.
The instance-level branch, on the other hand, focuses on training an instance-level classifier capable of distinguishing between positive and negative instances, despite the absence of explicit instance-level labels. To achieve this, it uses the attention scores $a_{i}^{j}$ from the bag-level branch as soft pseudo-labels, enabling weakly supervised learning at the instance level. The instance branch outputs a prediction $\hat{y}_{i}^{j}$ for each instance $x_{i}^{j}$. We note that $\hat{y}_{i}^{j}$ is passed through the hard instance mining mask and a max pooling layer, and then applied to the bag branch as a supervision signal in addition to the bag label.
Next, we discuss in details the architectures of the bag and instance branches of the model, as well as the design of the loss functions that leverages W2S generalization capability of the model.
3.4 Bag Branch
The training of the bag-level branch is conducted at the granularity of individual bags, meaning that each training batch consists of image patches that all originate from the same bag. By processing all patches belonging to a single bag simultaneously, the model is able to learn a feature aggregator which produces informative bag-level representations by compressing instance features.
In the bag-level branch, image patches $\{x_{i}^{j}\}_{j=1}^{n_{i}}$ within the bag $X_{i}$ are first passed through an image encoder $\mathcal{E}$ to extract their corresponding feature vectors $\{h_{i}^{j}\}_{j=1}^{n_{i}}$. These features are then processed by a hard positive mining module, which filters out the most informative instances. More specifically, given the vector of attention score $a_{i}$ for bag $X_{i}$, the hard positive mining module $\mathcal{M}$ sets the normalized attention score to 00 if the output of the instance classifier $\hat{y}_{i}^{j}$ exceeds a threshold:
$\tilde{\alpha}_{i}^{j}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathcal{M}\left(\alpha_{i}\right)^{j}=\left\{\begin{array}[]{ll}0&\text{if }\sigma\left(\hat{y}_{i}^{j}\right)\geq\tau\\ \alpha_{i}^{j}&\text{otherwise}\end{array}\right.$ (6)
where $\sigma(\cdot)$ is the sigmoid function. We further denote by $M_{i}$ the set of instances in $X_{i}$ with $\alpha_{i}^{j}>0$.
The features $\{h_{i}^{j}\}_{j=1}^{n_{i}}$ are aggregated via an attention-based aggregator $G$ to generate a global feature representation of the entire bag, denoted by $H_{i}=\sum_{j\in[n_{i}]}\tilde{\alpha}_{i}^{j}\cdot h_{i}^{j}$. Finally, the bag-level feature vector $H_{\text{bag}}$ is input to a bag-level classifier $\psi$ obtain predictions $\widehat{Y_{i}}$ for bag classification $\widehat{Y}_{i}=\psi(H_{i})$.
3.5 Instance Branch
The instance branch is used to force the shared encoder to learn representations for all instances, rather than only the instances with high attention scores in the bag branch. Model training for the instance branch uses instances from different bags in each batch, and this treatment allows the model to learn an overall distribution of positive instance features and a unified instance classifier across bags.
The instance branch consists of the encoder $\mathcal{E}$ shared with bag branch, and an instance classifier. The instances from different bags encoded to features $h_{i}^{j}$ are input to the instance classifier $\phi$ to obtain instance predictions $\hat{y}_{i}^{j}=\phi(h_{i}^{j})$. During training, the encoder $\mathcal{E}$ and the instance-level classifier $\phi$ are trained using weak supervision from the bag branch. Specifically, we pass the attention weights $a_{i}^{j}$ from the bag classifier through a sigmoid function, and use the output as soft labels for instances.
3.6 Loss Functions
The bag branch: The loss function for the bag-level branch comprises four components, collectively guiding the model’s optimization process in a comprehensive manner.
$L_{\text{bag}}=L_{\text{label}}+\beta\cdot L_{\text{inst}}+\gamma\cdot L_{\text{self}}+\delta\cdot L_{\text{attn}}\,,$ (7)
in which $\beta$, $\gamma$, and $\delta$ are hyperparameters. We postpone the discussion about the principle for selecting them.
Next we explain all the terms, and how they collectively correct pseudo label errors.
The first term $L_{\text{label}}$ is defined as
$L_{\text{label}}=c_{1}\cdot CE\left(\widehat{Y}_{i},Y_{i}\right)+c_{2}\cdot CE\left(\hat{y}_{i}^{k},Y_{i}\right)\,,$ (8)
where $k=\text{argmax}_{j\in M_{i}}(a_{i}^{j})$ is the instance with the largest attention score. Inspired by CLAM, $L_{\text{label}}$ is defined as a weighted sum of the cross-entropy between the bag label and two predictions: the bag-level prediction from the bag branch, and the instance-level prediction corresponding to the instance with the highest attention score.
During the early stages of training, $L_{\text{label}}$ plays a dominant role in guiding model updates and directly influences the subsequent convergence behavior of the model. Therefore, we assigned the highest weight among the loss components in the bag branch. This dual-component loss can captures both the global semantic understanding of the entire bag and the local discriminative power of key instances.
The second term $L_{\text{inst}}$ is defined as
$L_{\text{inst}}=CE\left(\widehat{Y}_{i},\hat{y}_{i}^{j}\right)\,,\text{ where }\hat{y}_{i}^{j}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max\limits_{j\in M_{i}}\left(\hat{y}_{i}^{j}\right)$ (9)
$L_{\text{inst}}$ is used to provide supervision signal from the instance branch. Specifically, the instance-level classifier predicts all patches in each bag and aggregates the filtered patch-level predictions by max-pooling operation.
This loss term serves to guide the instance-level classifier to identify and highlight the most discriminative patches in each bag, thus improving the final bag-level prediction. Ultimately, the goal of $L_{\text{inst}}$ is to help the bag-level classifier focus on key regions more accurately through indirect supervision, thus improving the overall classification performance.
The definition of $L_{\text{self}}$ is given as:
$L_{\text{self}}=c_{1}\cdot CE\left(\widehat{Y}_{i},I_{t}\left(\widehat{Y}_{i}\right)\right)+c_{2}\cdot CE\left(\hat{y}_{i}^{j},I_{t}\left(\hat{y}_{i}^{j}\right)\right)$ (10)
where $I_{t}(\cdot):\mathbb{R}\to\{0,1\}$ is an indicator function with a threshold $t$, i.e. $I_{t}(z)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{I}(z>t\,)$. It takes the value of $1$ if the confidence on the prediction reaches a certain level.
$L_{\text{self}}$ is a self-confidence loss function, it encourages the model to enhance well-learned features while preventing them from being easily altered. The loss utilizes the generalization capability of the bag branch by weighing on its own confident predictions.
Remark (A tripartite balance between $L_{\text{label}}$, $L_{\text{inst}}$, and $L_{\text{self}}$).
A bonus of adding $L_{\text{self}}$ to the bag loss is to form a tripartite balance between three sources of supervision: the bag label, the weak supervision from the instance model, and the self-generated hard label. Since the training is performed on $\{x_{i}^{j}\}_{j\in{M_{i}}}$, the original bag label is no longer a ground truth for the masked bag. During the model training process, some of the bags passing through the hard positive mining module may filter out all the positive instances, potentially introducing noise to their labels. On the other hand, the instance branch can only provide weak supervision for the bags. The W2S generalization capability is leveraged by $L_{\text{self}}$ to correct pseudo label errors when 1) the soft label is not confident; 2) the bag label and the instance branch supervision signal disagree.
The first three loss terms $L_{\text{label}}$, $L_{\text{inst}}$, and $L_{\text{self}}$ together determine the update of the bag classification head in each bag iteration. To ensure the tripartite balance, the weight of one loss should not exceed the sum of the weights of the other two.
We note that there is a potential mismatch in concepts by directly using normalized attention scores to supervise instance branch training. The score may attend to instances equally important for all classes, but not paying enough attention on discriminative features for distinct classes. To address the issue, we propose an additional loss $L_{\text{attn}}$ to calibrate the normalized attention scores.
$L_{\text{attn}}=CE(\max\limits_{j}\sigma(a_{i}^{j}),Y_{i})\,.$ (11)
We pass the highest attention score of all instances in a bag through a sigmoid function and align it with the bag label. It encouraging the model to learn a correlation between a bag class and the attention score of its most attended instance. Adding the loss effectively improves the quality of soft labels as weak supervision for the instance branch.
The instance branch: The loss function for the instance branch has two terms. For each instance $x_{i}^{j}$, we let $\ell_{\text{inst}}=\ell_{\text{pseudo}}+\theta\,\cdot\,\ell_{\text{self}}$ , where $\theta$ is a hyperparameter. The first term $\ell_{\text{pseudo}}$ is defined as $\ell_{\text{pseudo}}=CE(\hat{y}_{i}^{j},\sigma^{\prime}(a_{i}^{j}))$, where $\sigma^{\prime}(a_{i}^{j})=\sigma(a_{i}^{j})$ if $Y_{i}=1$ and $\sigma^{\prime}(a_{i}^{j})=0$ otherwise.
To enhance the robustness of the training process, a self-confidence loss term $\ell_{\text{self}}$ is introduced into the loss function to regulate the uncertainty associated with these pseudo-labels. It is defined as $\ell_{\text{self}}=CE(\hat{y}_{i}^{j},I_{t}(\hat{y}_{i}^{j}))$. $I_{t}(\cdot)$ is defined in the same way as in the bag branch.
3.7 Scheduling Dual-Level Training
The training algorithm improves the model on both bag- and instance- level prediction tasks, which resembles a multi-task learning setting. The bag training process updates parameters in the shared encoder, the attention module, and the bag prediction head. Its instance counterpart updates parameters in the shared encoder and the instance classification head.
Since the bag branch is more complex, we adopt a simple strategy to keep the optimization of both branches at roughly the same pace: we use a periodic schedule where in each cycle we first train the bag branch for $\kappa$ iterations, and then the instance branch for one iteration. One mini-batch is used for optimization in each iteration. More advanced scheduling algorithms are left for further study.
4 Experiments
4.1 Datasets
To comprehensively evaluate the performance of our method, we conducted experiments on two real-world datasets. CAMELYON16 [42] is a benchmark dataset for evaluating metastasis detection algorithms, consisting of $398$ WSIs of lymph node sections stained with hematoxylin and eosin (H&E), $269$ of which are in the training set and $129$ are in the test set. Besides slide labels, it also provides annotated contours indicating ground-truth patch labels. A patch is labeled positive if at least 25% of its area is annotated as lesion tissue. Instance labels are used exclusively for testing and remain inaccessible during training. We conduct experiments on the dataset under two settings: (1) using the original training and test sets directly for training and evaluation, and (2) creating a validation set by extracting 54 WSIs from the original training set, yielding a split of 215 WSIs for training, 54 for validation, and 129 for testing. We denote these two settings as C16 and C16v, respectively.
TCGA-NSCLC [^1] Official website for TCGA: https://www.cancer.gov/ccg/research/genome-sequencing/tcga has two cancer subtypes, lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC). The dataset contains $478$ LUAD slides and 478 LUSC slides, with only slide labels provided. All experiments on TCGA-NSCLC are conducted using $4$-fold cross-validation, with the data split into training, validation, and test sets in a ratio of 13:2:5.
4.2 Evaluation Metrics and Implementation Details
Evaluation Metrics: We report the area under the curve (AUC) and accuracy (ACC) of each method on all datasets. Both bag- and instance-level AUC and ACC are presented on CAMELYON16, however, only bag-level evaluation is available on TCGA-NSCLC due to the absence of ground-truth instance labels.
Implementation Details: Each WSI slide is cropped into a series of non-overlapping patches with a size of 512 $\times$ 512, and background regions are removed. A ResNet-50 network pretrained on the ImageNet is employed to extract a 1024-dimensional embedding vector from each patch for the training step. The shared encoder is a fully connected layer followed by a non-linear module. During training, we use Adam optimizer with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-5}$. The mini-batch sizes of the bag and the instance branch are $1$ and $1024$, respectively. The bag-level training and instance-level training are scheduled to run periodically. In each cycle, the algorithm run $10$ epochs of bag-level training, followed by one epoch of instance-level training. All models are trained for $200$ epochs on CAMELYON16 in the C16 setting, and at most $200$ epochs with early stopping in the C16v setting. For TCGA-NSCLC, early stopping is applied during training, with a maximum of $200$ epochs and patience set to $20$. All experiments were performed on a machine with 8 NVIDIA GeForce RTX 3090 GPUs.
4.3 Baselines
We compare our method with existing methods, including ABMIL, DSMIL, CLAM, and TransMIL. We also implement existing end-to-end boosting frameworks, WENO and MHIM-MIL. All boosting frameworks are tested by combining with ABMIL and DSMIL, respectively. We reproduced the baseline methods based on their published code [^2] A unified framework for MIL methods: https://github.com/lingxitong/MIL_BASELINE (under the MIT license); official implementation of WENO: https://github.com/miccaiif/WENO (unlicensed); official implementation of MHIM-MIL: https://github.com/DearCaat/MHIM-MIL (unlicensed)., with necessary modifications to adapt to different dataset settings. We exclude CIMIL and CausalMIL from the baselines since they have multi-stage training processes, which significantly increases computational and implementation complexity. For methods without instance-level prediction output, we find a reasonable approach to equip the model with such a capability and report the result.
4.4 Results
Table 1 shows the performance of binary classification between positive and negative on CAMELYON16 and subtyping between LUAD and LUSC on TCGA-NSCLC. Results of all models are presented, except for the fully supervised method on TCGA-NSCLC due to the lack of instance labels, and the instance-level results of TransMIL on CAMELYON16 since the features of instances are fused in several modules and therefore no instance prediction is available.
Table 1: Results on the CAMELYON16 and TCGA-NSCLC dataset
Methods C16 TCGA-NSCLC
Bag Instance Bag
AUC ACC AUC ACC AUC ACC
Fully supervised 0.8961 0.8915 0.9759 0.9584
MaxPooling 0.6128 0.7083 0.6857 0.7244 0.9289 0.8462
MeanPooling 0.6012 0.6927 0.8637 0.8746 0.8770 0.7848
ABMIL 0.8508 0.8372 0.8067 0.7341 0.9261 0.8618
DSMIL 0.7392 0.7674 0.8913 0.7779 0.8581 0.7904
CLAM-SB 0.8253 0.8450 0.8124 0.7402 0.9308 0.8646
CLAM-MB 0.8270 0.8605 0.8202 0.7566 0.9317 0.8593
TransMIL 0.9332 0.8372 0.9370 0.8450
ABMIL+WENO 0.9231 0.9070 0.8147 0.9120 0.9418 0.8996
DSMIL+WENO 0.9207 0.8966 0.9139 0.9282 0.9421 0.9011
ABMIL+MHIM-MIL 0.8969 0.8880 0.7895 0.6679 0.9504 0.9043
DSMIL+MHIM-MIL 0.9094 0.8906 0.8346 0.7146 0.9560 0.9154
ABMIL+Ours 0.9318 0.9173 0.9216 0.9291 0.9626 0.9121
DSMIL+Ours 0.9222 0.8915 0.9298 0.9288 0.9614 0.9077
CLAM-SB+Ours 0.9131 0.8915 0.9628 0.9334 0.9521 0.8912
CLAM-MB+Ours 0.9109 0.8915 0.9549 0.9355 0.9405 0.8804
From the results, MaxPooling and MeanPooling perform poorly on CAMELYON16, primarily due to their inability to effectively identify key instances within positive bags, tending to be disturbed by abundant negative instances. Attention-based methods, including ABMIL, DSMIL, CLAM, and TransMIL, utilize the attention mechanism to determine the significance of each patch and therefore avoid indiscriminately aggregating all instance information, resulting in higher performance over the pooling methods. However, their generalization ability is still limited due to dependence on a few instances with high attention scores seen during training.
Table 2: Results (AUC) of ABMIL and DSMIL with boosting frameworks
Methods C16 TCGA-NSCLC
ABMIL 0.8508 0.9261
+WENO 0.9231 (+0.0723) 0.9418 (+0.0157)
+MHIM-MIL 0.8969 (+0.0461) 0.9504 (+0.0243)
+Ours 0.9318 (+0.0810) 0.9626 (+0.0365)
DSMIL 0.7392 0.8581
+WENO 0.9207 (+0.1815) 0.9421 (+0.0840)
+MHIM-MIL 0.9094 (+0.1702) 0.9560 (+0.0979)
+Ours 0.9222 (+0.1830) 0.9614 (+0.1033)
Table 3: Results on CAMELYON16 (under the C16v setting)
Methods Bag Instance
AUC ACC AUC ACC
TransMIL 0.8546 0.8047
ABMIL+WENO 0.7703 0.8125 0.8579 0.9162
DSMIL+WENO 0.8245 0.8516 0.9087 0.9069
ABMIL+Ours 0.8553 0.8750 0.9117 0.9149
DSMIL+Ours 0.8602 0.8828 0.9102 0.9022
With augmented supervision in both bag and instance branches, our method shows remarkable classification ability on both bag and instance levels. As shown in Table 1, the combinations of our framework and the four attention-based MIL methods (i.e., ABMIL, DSMIL, CLAM-SB, and CLAM-MB) outperform the fully supervised method on the bag level. Also, CLAM-SB+Ours and CLAM-MB+Ours achieve the best performance on the instance level among all weakly supervised models. In addition, Table 3 gives a clearer comparison between our method and the competing boosting frameworks, WENO and MHIM-MIL, showing that our method significantly enhances the performance of ABMIL and DSMIL on both datasets, superior to the other two frameworks. Table  1 and 3 jointly demonstrate that our method (adapted to ABMIL and DSMIL) reaches the highest AUC under the C16v setting and on TCGA-NSCLC dataset.
Visualization of the Instance Predictions on the CAMELYON16 Dataset: In Figure 2 we compare the instance prediction results of our method combined with DSMIL and CLAM-SB, and results of WENO combined with DSMIL. The heatmaps clearly show that our model is more discriminative since it shows a greater contrast between the metastatic lesion area and the normal tissue.
(a) Ground-truth
(b) DSMIL+WENO
(c) DSMIL+Ours
(d) CLAM-SB+Ours
Figure 2: Visualization of instance-level prediction results. (a) shows one original CAMELYON16 slide thumbnail with ground-truth positive instances masked in green. The following are heatmaps overlapped on the slide image where deeper red indicates higher predicted probability of being positive and deeper blue indicates lower negative probability.
Ablation Study: To examine the effect of adding self-confidence loss to both the bag branch and the instance branch, we conduct ablation study on the CAMELYON16 dataset (in the C16 setting). We use the ABMIL model in a modified version of our method where $L_{\text{self}}$ and $\ell_{\text{self}}$ are removed from the loss functions. The bag-level (instance-level) AUC drops from 0.9318 (0.9216) to 0.9151 (0.9117), with a reduction of 0.0167 (0.0099). The bag-level (instance-level) ACC decreased from 0.9173 (0.9291) to 0.8915 (0.9149), with a drop of 0.0258 (0.0142).
5 Conclusion
In this paper, we propose a bi-directional distillation framework to enhance dual-level multiple instance learning by denoising weak supervision signals through mutual knowledge distillation. By aligning the soft labels of each branch with the predictions of its counterpart, the framework enables the bag-level and instance-level modules to collaboratively refine data representations and improve discriminative performance at both levels. By introducing self-confidence losses to both branches, the model further leverages the pseudo-label correction capability of W2S generalization to boost its prediction performance. Experiments on the CAMELYON16 and TCGA-NSCLC datasets show that our method consistently outperforms previous distillation-based boosting methods, especially for instance predictions.
Limitations of our work include the absence of theoretical guarantees and a lack of discussion on the fundamental capability limits of the mutual-improvement paradigm. MIL has been applied to WSI-based diagnosis in real-world settings. There exist potential negative societal impacts such as mistrust in AI-generated predictions.